OpenStudy (anonymous):

Simplify to lowest terms a^2 b^2 _____minus_____ a+b a+b is the answer ab^2

OpenStudy (anonymous):

Is my answer correct.....I am so confused

OpenStudy (saifoo.khan):

I just emailed you the solution! :)

OpenStudy (anonymous):

I did not get it right. I think I am on the right track with the next one. I know you cancel out. Right? You taught me that.

OpenStudy (anonymous):

\[\frac{ a^2 }{ a+b } - \frac{ b^2 }{a+b} \] Now since they have the same denominator, You'll simply write it this way, \[\frac{ a^2 - b^2 }{ a+b } \]. However the numerator has yet to be factorized, so.. \[\frac{ (a-b) (a+b) }{ a+b}\] Now cancel (a+b) from numerator with the same one from denominator..Eventually you'll get ( a-b ) as your final answer. :)